# SoK: Runtime Integrity

### v2

@@ 4 Effects & Consequences
This section discusses effects and consequences of design choices discussed in Section 3.
@@ 4.1.1 Coverage
In terms of coverage, CFI mechanisms offer varying degrees of protection, ranging from protecting all edges, i.e., all types of indirect control flow altering instructions [15, 86, 87, 88, 89, 123, 105, 124], to partial coverage, targeting either forward-edges [19, 118, 96, 107, 108, 110, 125, 23, 98] or backward-edges [126, 97, 28, 109, 127, 128, 21, 22]. Forward-edge [10] schemes and employs software-based IRMs [19, 118, 125], hardware-assisted monitoring [81, 105], landing pads [96, 23], and pointer authentication [98, 27, 90]. Backward edge [9] schemes utilize software-[129, 127] or hardware-based [22, 21] shadow stacks, and architecture-specific features, such as branch history tables in the x86 processors [130, 131], static rewriting, e.g., to form jump tables [109], or pointer authentication [97, 28].
@@
The majority of CFI mechanisms (regardless of their coverage) is context insensitive [15, 19, 89, 90, 87, 133, 86, 118, 23, 124]. This can introduce gaps that are challenging to detect [64]. Moreover, it limits the ability to detect non-control-data attacks [11].
@@ 4.1.2 Compatibility
CFA designs typically do not require source code knowledge to generate control flow evidence [34], whereas source code knowledge may assist in analyzing received evidence (see Section 5.2). Hardware-based CFA inherently supports binaries [36, 85, 72, 38] by integrating with the CPU core and detecting branch instructions at runtime. CFA relying on instrumentation [34, 71, 35, 84, 136] can instrument control flow transfers in the binary without knowing the source. This is because required instrumentation simply used only to log destination addresses, rather than determining/enforcing policies in place. Exceptions to this include schemes mixing evidence generation with local integrity checks, e.g., [120].
Modular/Shared Object Support. A limitation of many CFI mechanisms is the lack of support for external modules or dynamic shared objects (DSO). These mechanisms often rely on global information that may not always be available, making it challenging to implement globally compatible CFI. Abstractly speaking, support for external/shared modules involves (i) integrating multiple modules hardened by CFI separately and (ii) integrating CFI-protected modulfes with unprotected legacy code. Binary solutions such as CCFIR [87] attempt to address these issues by allowing more targets than necessary, striking a security-compatibility compromise. Although approaches such as MCFI [89] and RockJIT [137] tackle case (i) by independently instrumenting each module and generating new CFGs when modules are linked, recent CFI solutions that offer stronger security guarantees, exemplified by CFI [81] and OS-CFI [107], do not provide modular support. Even contemporary solutions employing hardware features, e.g., PACStack [28] and PACTight [98], struggle to address both issues (i) and (ii).
@@
CFA mechanisms that instrument binaries may also encounter compilation failures due to instrumentation issues, as observed in ReCFA with specific benchmarks [121]. As CFA is a newer concept, fewer studies exist on analyzing CFA instrumentation compatibility. At least in principle, issues presented in CFI schemes (e.g., similar to VIP-CFI, TinyCFA employs a reserved register) could also apply to CFA, depending on the instrumentation strategy used.
@@ 4.1.3 Feasibility
CFI approaches are to a large extent feasible, despite inherent uncertainties around the robustness of policies due to granularity and context sensitivity (discussed further in Section 4.2).
As with any attestation mechanism, CFA requires a secure RoT on to maintain and authenticate evidence, as discussed in Section 3.2. It also requires communication with an external . Naturally, custom hardware features (such as branch monitors) improve feasibility and reduce the cost of CFA. Being a relatively recent concept, we expect hardware features to support CFA to take longer to reach off-the-shelf devices.
@@ 4.1.4 Performance & Scalability
In contrast, CFA faces further scalability challenges in storing and transmitting runtime evidence. Schemes such as ScaRR [117] and ACFA [38], which continuously report evidence to , may face challenges when attempting to cover multiple active applications on the same . This may impact availability, particularly in scenarios where network communication is essential (e.g., cloud). In other schemes, may need to store a large CFLog is attested operations are complex. This can limit applicability to small and self-contained operations [71, 35].
@@ 4.2.1 Pitfalls
Corner Cases: Certain attacks exploit exceptional cases. For instance, the Control Jujutsu attack [145] highlighted the limitations of fine-grained CFI defenses with activated shadow stacks in complex code bases like Apache and nginx. Due to coding practices in these code bases, context-insensitive analysis regardless of its intended robustness creates over-approximated CFGs that render CFI ineffective. Another example is CHOP [146], which further undermines robust backward edge protection mechanisms, including hardware-based shadow stack implementations [22]. It leverages a specific corner case that enables manipulation of the stack unwinding path during exception handling to launch ROP-like attacks, using the unwinder as a confused deputy.
@@ 4.2.2 Control Flow Bending (CFB)
CFB attacks [20] generalize non-control data attacks targeting CFI schemes relying on statically generated CFGs. While many CFI attacks targeted weaker or sub-optimal implementations [101, 100, 99], CFB focuses on bypassing the most restrictive (or optimal) static CFI policies. CFB creates malicious (Turing-complete) paths that exist on the most strict CFG for a given program by exploiting specific functions, called dispatchers, which have the capability to modify their own return addresses. In other words, CFB can arbitrarily modify (bend) a program’s behavior/path while staying within the confines of the imposed security policy.

### v3

@@ 4 Effects & Consequences
This section discusses the effects and consequences of design choices presented in Section 3.
@@ 4.1.1 Coverage
In terms of coverage, CFI mechanisms offer varying degrees of protection, ranging from protecting all edges, i.e., all types of indirect control flow altering instructions [14, 88, 89, 90, 91, 127, 107, 128], to partial coverage, targeting either forward-edges [18, 121, 98, 109, 110, 112, 129, 22, 100] or backward-edges [130, 99, 27, 111, 122, 131, 20, 21]. Forward-edge [9] schemes employ IRMs via SWI [18, 121, 129], hardware-assisted monitoring [84, 107], landing pads [98, 22], and pointer authentication [100, 26, 92]. Backward edge [8] schemes utilize software-[132, 122] or hardware-based [21, 20] shadow stacks, and architecture-specific features, such as branch history tables in the x86 processors [133, 134], static rewriting, e.g., to form jump tables [111], or pointer authentication [99, 27].
@@
The majority of CFI mechanisms (regardless of their coverage) are context insensitive [14, 18, 91, 92, 89, 136, 88, 121, 22, 128]. This can introduce gaps that are challenging to detect [69]. Moreover, it limits the ability to detect non-control-data attacks [10].
@@
Device Type/Target | Mechanism | Scope | Overheads | ||||||||||||||
Year | Scheme |
Embedded (bare-metal) |
Embedded (OS) |
High-end (User-space) |
High-end (Kernel) |
Type |
Strategy |
Sensitivity | System Support |
Data-Only |
ROP |
JOP |
Evidence Expressiveness |
Runtime |
Code Size |
Custom Hardware |
Network |
Control Flow Integrity (CFI) Approaches | |||||||||||||||||
2013 | bin-CFI [88] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
2013 | CCFIR [89] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
2014 | LLVM CFI [18] | ✗ | ✗ | ✓ | ✓ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2014 | KCoFI [90] | ✗ | ✗ | ✗ | ✓ | Enforcement | SWI | ✗ | MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2014 | MCFI [91] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | CS | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2014 | RockJIT [137] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | CS | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2015 | CCFI [92] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2015 | HAFIX [130] | ✓ | ✗ | ✗ | ✗ | Enforcement | ISA | Path | C-HW | ✗ | ● | ✗ | - | ⚫ | ⚫ | ⚫ | - |
2015 | CFCI [93] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
2015 | O-CFI [94] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | ✗ | OS/MMU+MPX | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2015 | CFI [85] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2015 | PathArmor [106] | ✗ | ✗ | ✓ | ✗ | Hybrid | SWI+ISA | Path | OS/MMU+LBR | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2015 | Lockdown [95] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | ✗ | OS/MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2016 | TypeArmor [96] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2016 | FG-CFI [97] | ✗ | ✗ | ✗ | ✓ | Enforcement | SWI | ✗ | MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2016 | HCFI [127] | ✗ | ✗ | ✓ | ✗ | Enforcement | ISA | ✗ | OS+C-HW | ✗ | ● | ● | - | ⚫ | ⚫ | ⚫ | - |
2017 | PittyPAT [107] | ✗ | ✗ | ✓ | ✗ | Hybrid | ISA+HRM | Path | OS/MMU+PT | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2017 | GRIFFIN [128] | ✗ | ✗ | ✓ | ✗ | Hybrid | ISA+HRM | ✗ | OS/MMU+PT+TSX | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2017 | CFI-CaRE [136] | ✓ | ✗ | ✗ | ✗ | Enforcement | SWI+R/I | ✗ | TZ | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
2017 | Intel CET [21] | ✗ | ✗ | ✓ | ✓ | Enforcement | ISA | ✗ | OS/MMU+CET | ✗ | ● | ○ | - | ⚫ | ⚫ | ✗ | - |
2018 | CFI [84] | ✗ | ✗ | ✓ | ✗ | Hybrid | SWI+ISA | ✗ | OS/MMU+PT | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2018 | SCFP [108] | ✓ | ✗ | ✗ | ✗ | Enforcement | SWI+C-HW | Path | C-HW | ✗ | ● | ◐ | - | ⚫ | ⚫ | ⚫ | - |
2018 | ARM BTI [98] | ✓ | ✓ | ✓ | ✓ | Enforcement | ISA | ✗ | BTI | ✗ | ✗ | ○ | - | ⚫ | ⚫ | ✗ | - |
2018 | PAC-RET [99] | ✓ | ✓ | ✓ | ✓ | Enforcement | ISA | ✗ | PA | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
2019 | OS-CFI [109] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | Path | OS/MMU+MPX+TSX | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
2019 | CFI-LB [110] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | CS | OS/MMU+TSX | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
2019 | PARTS [25] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+ISA | ✗ | OS/MMU+PA | + | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2020 | RAI [111] | ✓ | ✗ | ✗ | ✗ | Enforcement | SWI+R/I | Path | MPU | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
2020 | Silhouette [122] | ✓ | ✗ | ✗ | ✗ | Enforcement | SWI+R/I | ✗ | MPU | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
2021 | VIP [112] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+R/I | Path | OS/MMU+MPK | + | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
2021 | PACStack [27] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+ISA | ✗ | OS/MMU+PA | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
2022 | TyPro [129] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | ✗ | OS/MMU | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
2022 | PAL [26] | ✗ | ✗ | ✗ | ✓ | Enforcement | SWI+ISA | ✗ | PA+MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2022 | PACTight [100] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI+ISA | ✗ | OS/MMU+PA | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
2023 | FineIBT [22] | ✗ | ✗ | ✓ | ✓ | Enforcement | SWI+ISA | ✗ | CET+MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
2023 | SHERLOC [114] | ✓ | ✗ | ✗ | ✗ | Hybrid | ISA+HRM | Path | TZ+MTB+DWT | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
2023 | TypeSqueezer [138] | ✗ | ✗ | ✓ | ✗ | Enforcement | SWI | Path | OS/MMU | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
2024 | HEK-CFI [139] | ✗ | ✗ | ✗ | ✓ | Enforcement | ISA | ✗ | CET+MMU | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
Control Flow Attestation (CFA) Approaches | |||||||||||||||||
2016 | C-FLAT [33] | ✓ | ✗ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2017 | LO-FAT [35] | ✓ | ✗ | ✗ | ✗ | Monitoring | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
2017 | ATRIUM [36] | ✓ | ✗ | ✗ | ✗ | Monitoring | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
2018 | LiteHAX [37] | ✓ | ✗ | ✗ | ✗ | Monitoring | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
2019 | DIAT [140] | ✓ | ✓ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2019 | ScaRR [119] | ✗ | ✗ | ✓ | ✗ | Monitoring | SWI | -based | OS/MMU | ● | ● | ⚫ | ⚫ | ✗ | |||
2019 | RIM [124] | ✓ | ✗ | ✗ | ✗ | Monitoring | C-HW | Path | C-HW | ● | ● | ✗ | ✗ | ? | |||
2020 | OAT [38] | ✓ | ✓ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2020 | LAHEL [141] | ✓ | ✓ | ✗ | ✗ | Monitoring | C-HW | -based | C-HW | ◐ | ◐ | ⚫ | ⚫ | debug HW | |||
2020 | LAPE [142] | ✓ | ✗ | ✗ | ✗ | Monitoring | SWI+R/I | -based | MPU | ◐ | ◐ | ⚫ | ⚫ | ✗ | |||
2021 | Tiny-CFA [34] | ✓ | ✗ | ✗ | ✗ | Monitoring | SWI | -based | C-HW | ● | ● | ⚫ | ⚫ | ✗ | |||
2021 | DIALED [76] | ✓ | ✗ | ✗ | ✗ | Monitoring | SWI | -based | C-HW | ● | ● | ⚫ | ⚫ | ✗ | |||
2021 | ReCFA [125] | ✗ | ✗ | ✓ | ✗ | Monitoring | SWI+R/I | -based | OS+MPK | ● | ● | ⚫ | ⚫ | ✗ | |||
2022 | GuaranTEE [126] | ✗ | ✗ | ✓ | ✗ | Monitoring | SWI | -based | Intel SGX | ● | ● | ⚫ | ⚫ | ✗ | |||
2023 | ACFA [41] | ✓ | ✗ | ✗ | ✗ | Monitoring | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
2023 | ARI [40] | ✓ | ✓ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ◐ | ◐ | ⚫ | ⚫ | ✗ | |||
2023 | BLAST [39] | ✓ | ✓ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2023 | ISC-FLAT [43] | ✓ | ✗ | ✗ | ✗ | Hybrid | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2024 | TRACES [42] | ✓ | ✗ | ✗ | ✗ | Monitoring | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
2024 | CFA+ [120] | ✓ | ✓ | ✓ | ✗ | Hybrid | SWI+ISA | -based | OS/MMU+TPM | ● | ● | ⚫ | ⚫ | ✗ |
Legend: ✓ Has this feature, ✗ Lacks this feature, - Feature is not applicable, SWI: Software Instrumentation, HRM: Hardware Reference Monitor, R/I: Randomization or Isolation, ISA: Instruction Set Architecture, MMU: Memory Management Unit, C-HW: Custom Hardware, OS: Operating System, CS: Context Sensitive, Path: Context- & path-sensitive MPX: Intel Memory Protection eXtensions LBR: Intel Last Branch Record PT: Intel Processor Trace TSX: Intel Transactional Synchronization Extension TZ: ARM TrustZone, CET: Intel Control-Flow Enforcement Tech., BTI: ARM Branch Target Identification, PA: ARM Pointer Authentication, MPU: Memory Protection Unit, MPK: Intel Memory Protection keys, MTB: ARM Macro Trace Buffer, DWT: ARM Data Warchpoint and Trace, Definition-to-use data corruption, Path deviation DOP, Definition-to-use data corruption & path deviation DOP, ● Fine-grained, ◐ Mixed granularity, ○ Coarse-grained, No path or hashed path, Partial path, Full path (lossless), All evidence stored and transmitted at once, Evidence sliced and streamed, ✗ No overhead, ⚫ 5% overhead, ⚫ 20% overhead, ⚫ Higher overhead, ? Feature required but cost not reported, debug HW reliance on prototyping/debug features not meant for device deployment.
@@ 4.1.2 Compatibility
CFA designs typically do not require source code knowledge to generate control flow evidence [33], whereas source code knowledge may assist in analyzing received evidence (see Section 5.2). Hardware-based CFA inherently supports binaries [35, 36, 37, 41] by integrating with the CPU core and detecting branch instructions at runtime. CFA relying on SWI [33, 38, 34, 43, 39, 42] can instrument control flow transfers in the binary without knowing the source. This is because required instrumentation is used only to log destination addresses rather than determining/enforcing policies in place. Exceptions to this include schemes mixing evidence generation with local integrity checks, e.g., [124, 120].
Modular/Shared Object Support. A limitation of many CFI mechanisms is the lack of support for external modules or dynamic shared objects (DSO). These mechanisms often rely on global information that may not always be available, making it challenging to implement globally compatible CFI. Abstractly speaking, support for external/shared modules involves (i) integrating multiple modules hardened by CFI separately and (ii) integrating CFI-protected modules with unprotected legacy code. Binary solutions such as CCFIR [89] attempt to address these issues by allowing more targets than necessary, striking a security-compatibility compromise. Although approaches such as MCFI [91] and RockJIT [137] tackle case (i) by independently instrumenting each module and generating new CFGs when modules are linked, recent CFI solutions that offer stronger security guarantees, exemplified by CFI [84] and OS-CFI [109], do not provide modular support. Even contemporary solutions employing hardware features, e.g., PACStack [27] and PACTight [100], struggle to address both issues (i) and (ii).
@@
CFA mechanisms that instrument binaries may also encounter compilation failures due to instrumentation issues, as observed in ReCFA with specific benchmarks [125]. As CFA is a newer concept, fewer studies exist on analyzing CFA instrumentation compatibility. At least in principle, issues presented in CFI schemes could also apply to CFA, depending on the instrumentation strategy used. (For instance, similar to VIP-CFI, TinyCFA and TRACES also employ reserved registers.)
@@ 4.1.3 Feasibility
CFI approaches are to a large extent feasible, despite inherent uncertainties around the robustness of policies due to granularity and context sensitivity (discussed further in Section 4.2). As with any attestation mechanism, CFA requires a secure RoT on to maintain and authenticate evidence, as discussed in Section 3.2. It also requires communication with an external . Naturally, custom hardware features (such as branch monitors) improve feasibility and reduce the cost of CFA. Being a relatively recent concept, we expect hardware features to support CFA to take longer to reach off-the-shelf devices.
@@ 4.1.4 Performance & Scalability
Performance and scalability in CFA depend on the monitoring strategy. Techniques that utilize SWI incur runtime and code-size increases due to logging [34], context switching into secure environments [39], or both [33, 140, 119, 38, 142, 40, 42]. Alternatively, hardware-based CFA mechanisms trade the performance/scalability impact of SWI for additional hardware costs. Early hardware-based approaches impose expensive overheads due to large internal buffers and hardware hash-engines [36, 35]. Alternative techniques leverage minimal hardware [37, 34] or propose hardware/software co-designs [41] to lower the hardware cost. Regardless of their specificities, existing hardware-based CFA mechanisms are tailored to embedded platforms, leaving scalability in more complex computing platforms as an open challenge.
Unlike CFI, CFA faces further performance challenges in storing and transmitting runtime evidence. Schemes such as ScaRR [119], ACFA [41], and TRACES [42], which continuously report evidence to , may face challenges when attempting to cover multiple active applications on the same . This may impact availability, particularly when network communication is essential (e.g., cloud). In other schemes, may need to store a large CFLog if attested operations are complex. The latter can limit CFA applicability to small and self-contained operations [38, 34]. Recent work has investigated dynamically configurable and application-specific optimizations for CFA evidence in MCUs [148]. Nevertheless, additional research is required to realize CFA (and related optimizations) in higher-end systems.
@@ 4.2.1 Pitfalls
Corner Cases: Certain attacks exploit exceptional cases. For instance, the Control Jujutsu attack [153] highlighted the limitations of fine-grained CFI defenses with activated shadow stacks in complex code bases like Apache and nginx. Due to coding practices in these code bases, context-insensitive analysis, regardless of its intended robustness, creates over-approximated CFGs that render CFI ineffective. Another example is CHOP [154], which further undermines robust backward edge protection mechanisms, including hardware-based shadow stack implementations [21]. It leverages a specific corner case that enables manipulation of the stack unwinding path during exception handling to launch ROP-like attacks, using the unwinder as a confused deputy.
@@ 4.2.2 Control Flow Bending (CFB)
CFB attacks [19] generalize non-control data attacks targeting CFI schemes relying on statically generated CFGs. While many CFI attacks target weaker or sub-optimal implementations [103, 102, 101], CFB focuses on bypassing the most restrictive (or optimal) static CFI policies. CFB creates malicious (Turing-complete) paths that exist on the most strict CFG for a given program by exploiting specific functions, called dispatchers, which have the capability to modify their own return addresses. In other words, CFB can arbitrarily modify (bend) a program’s behavior/path while staying within the confines of the imposed security policy.